OpenStudy (anonymous):

Anyone know how do this one The sum of two numbers is 33 and their product is 266. Find two number.

OpenStudy (anonymous):

yes, guess and check

OpenStudy (anonymous):

for example \(20\times 13=260\) which is too low, try \(19\times 14\)

OpenStudy (jennychan12):

system of equations. x+y = 33 and xy = 266 solve for x and plug it in. :)

OpenStudy (anonymous):

yup, got it on the second try

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

you just have to play around and see what works pick two numbers that add to 33 and see i picked 20 and 13 because it was easy to see that 20 + 13 = 33 but the product wasn't right, so i tried another to get a larger product

OpenStudy (anonymous):

 should it be something like this